the volume of air bubble at the bottom of a lake when the pressure is (3p/4)atm is (3v/8)cm³,what will be the product of the volume of the air bubbles at the surface of the lake of the pressure is 2/3 of the original value and the pressure is reduced by 3/8 of the original value at constant temperature?


1
Expert's answer
2021-01-19T04:36:36-0500

The question is two in one

I. P1= 3p/4

V1= 3v/8

P2= 2/3P1 = 2/3 x 3p/4 = p/2

V2= ?

From boyle's law

P1V1 = P2V2

V2 = P1V1/P2 = 3p/4 x 3v/8 ÷ p/2

= 9v/16


II. If P2= 3/8P1 = 3/8 x 3p/4 = 9p/32

V2= 3p/4 x 3v/8 x 32/9p = v